OpenStudy (anonymous):

how do u know if sin^-1(x) is (sin(x))^-1 or 1/sinx = csc(x)?

OpenStudy (helder_edwin):

the -1 power indicates the inverse function. as far as i know, only american books use this confusing notation. it is always better to write \(\arcsin x\) for the inverse of \(\sin x\).

OpenStudy (sirm3d):

the exponent (-1) is reserved for the inverse of a function. sin^-1(x) is always the arcsin x (the inverse of sin x). if the reciprocal is intended, then it will be written \[\huge \left( \sin x \right)^{-1}\]

OpenStudy (anonymous):

thanks. whats if its sin^-2(x)? is sin^10(x)=(sin(x))^10?

OpenStudy (helder_edwin):

good point. it is a convention: \[ \large \sin^nx=(\sin x)^n \] ergo the confution.

OpenStudy (sirm3d):

\[\huge \sin^{-2}x=\frac{ 1 }{ \sin^2 x }\]. such notation sin^(-n) x is avoided because of the different interpretations it brings

OpenStudy (anonymous):

so only true for n

OpenStudy (anonymous):